# EEG-Language Modeling for Pathology Detection

### v2

@@ Abstract
Multimodal language modeling constitutes a recent breakthrough which leverages advances in large language models to pretrain capable multimodal models. The integration of natural language during pretraining has been shown to significantly improve learned representations, particularly in computer vision. However, the efficacy of multimodal language modeling in the realm of functional brain data, specifically for advancing pathology detection, remains unexplored. This study pioneers EEG-language models trained on clinical reports and 15000 EEGs. We extend methods for multimodal alignment to this novel domain and investigate which textual information in reports is useful for training EEG-language models. Our results indicate that models learn richer representations from being exposed to a variety of report segments, including the patient’s clinical history, description of the EEG, and the physician’s interpretation. Compared to models exposed to narrower clinical text information, we find such models to retrieve EEGs based on clinical reports (and vice versa) with substantially higher accuracy. Yet, this is only observed when using a contrastive learning approach. Particularly in regimes with few annotations, we observe that representations of EEG-language models can significantly improve pathology detection compared to those of EEG-only models, as demonstrated by both zero-shot classification and linear probes. In sum, these results highlight the potential of integrating brain activity data with clinical text, suggesting that EEG-language models represent significant progress for clinical applications.
@@ 1 Introduction
Medical neuroimaging such as electroencephalography (EEG) has not yet benefited to the same extent as other domains from the considerable advances deep learning has brought about. While EEG sees widespread clinical use for pathology detection, in particular for epilepsy [1, 2] as well as sleep disorders [3], available annotated data is scarce. As the impressive scaling properties of deep learning are now well described [4, 5], self-supervised learning (SSL) is a promising direction by enabling pretraining with unlabeled data and thereby increasing available training sample sizes [6, 7]. Various such approaches have shown initial success when applied to EEG. These include methods relying on data-augmentations [8, 9], the temporal ordering of EEG data [10], as well as masking and reconstruction [11]. However, these are hindered by the difficulty of creating appropriate data augmentations and, especially reconstruction techniques, by low signal-to-noise. Thus, progress in the medical context has lagged, likely further exacerbated by the modality displaying high similarity between pathologies.
Figure 1: Overview of the methodology. A) The EEG and text encoding pathways are displayed. Arrows in orange (blue) denote operations only used by ELMe,l (ELMl). $\mathbf{h}_{e}$ and $\mathbf{e}$: EEG embedding prior and post projection, $\mathbf{h}_{l}$ and $\mathbf{l}$: text embedding prior and post projection B) Summary of the downstream tasks of EEG or clinical report retrieval and pathology detection. C) Visualization of the current extension of multimodal language modeling to EEG timeseries, which are cut into sequential crops.
@@
However, language-EEG pretraining also entails unique challenges. First, datasets are generally smaller than those used in radiology. As a consequence, models used for EEG tend to be smaller, which, when jointly trained with a much larger language model, may lead to unstable training and dimensional collapse of the latent space [16, 17]. Second, the clinical reports tend to be highly heterogeneous in both content and length. While previous applications have paired natural and medical images with short captions [12, 13], EEG reports tend to span multiple paragraphs and include information irrelevant to downstream clinical tasks, potentially hindering the pretraining process.
The current work investigates the application of aligning functional brain data with medical textual information for the first time. To address challenges associated with data scarcity, we integrate recent methodological developments in multimodal pretraining for medical imaging to regularize the latent space [17] and compare these to a contrastive learning approach. Furthermore, we investigate how to best handle the heterogeneity of medical EEG reports. Specifically, we perform content-based text segmentation, enabling inference on the relative importance of the different sources of information in the reports, such as clinical history, medication, description of the recording, and the physician’s impression. These approaches are additionally compared to the use of a large generative language model to summarize clinical reports, which may be steered for specific downstream tasks. To study the information learned by EEG-language models, we perform retrieval analyses of clinical reports or EEG. We furthermore test downstream performance on classifying normal and pathologically abnormal EEG, which is a widespread clinical task. These tests include zero-shot classification by leveraging the language capabilities to evaluate the flexibility of the approach. Consequently, we provide insight into the novel extension of current multimodal language modeling to the domain of medical brain activity. Our results constitute considerable increases in pathology detection performance in scenarios with few labels. These are particularly relevant for clinical contexts, which tend to operate with smaller datasets compared to many common areas of deep learning applications.
@@ 2 Related work
  • Self-supervised learning with EEG data. SSL with EEG data has been predominantly applied to emotion recognition [18, 19], motor imagery [20, 21], sleep staging [9, 21], as well as pathology detection. For the latter application, the temporal order of EEG crops was used initially to demonstrate label-efficient representation learning [10]. Contrastive learning, combined with larger EEG encoders trained on multiple datasets, further improved pathology detection [8]. In this case, the model was trained for invariance to a set of data augmentations, based on the SimCLR work in computer vision [7]. Whereas most work uses convolutional and recurrent architectures, recent studies have explored transformers [22, 11]. These have focused on scaling and adopted tokenization in an attempt to improve the challenge of effective cross-dataset EEG training. They rely on masking and reconstruction pretext tasks.
  • Using EEG for pathology detection. While SSL shows good performance for pathology detection, it is particularly in contexts with little annotated data that it performs well. When more labeled data is available, expert-based feature extraction combined with traditional machine learning classifiers are competitive together with supervised deep learning [23, 24, 25, 26, 27]. The observation of similar classification performance between deep learning and feature-based approaches has also been made for other EEG applications [28, 29]. As a consequence, it has been posited that label noise could induce a ceiling effect on classification performance [30, 24]; specifically, the inter-rater reliability of EEG classification into normal or pathologically abnormal by neurologists. Although still only a hypothesis, if this is so, a focus on improving low-label classification performance may be of extra importance.
  • Medical multimodal language modeling. Medical vision-language modeling aims to guide self-supervised pretraining on medical images using textual information in reports, with performance on a variety of downstream tasks benefiting as a result [31, 32, 13]. Recent results indicated that the finetuning of large language encoders is not necessary [17]. Due to less available data in the medical domain, using a pretrained encoder and freezing its weights considerably reduces computational cost while boosting downstream performance [17]. This was attributed to avoiding collapse in latent space by directly aligning image embeddings to the frozen text embeddings and additional regularization. Nevertheless, this line of work focuses mainly on the X-ray, CT images, and structural MRI images [33, 34]. Recently, multimodal language modeling has been performed using electrocardiogram data for cardiac pathology detection. Besides confirming the good performance using frozen language models, this work suggests effectiveness of combining functional timeseries data with textual information in a medical context [35, 36, 37, 38].
@@ 3.1.1 EEG-language pretraining
Here we detail the setup for pretraining EEG-language models (ELMs). Whereas vision-language models are typically trained by aligning a 2D image with a short caption [13, 12], EEG-language modeling is confronted with long EEG time series and multi-paragraph medical reports. To overcome this, we employ text segmentation and time series cropping to create multiple non-overlapping samples per modality and subject. Next, we propose sub-unit alignment by pretraining on these cropped samples. In addition to considerably increasing sample size, this enables the extension of successful approaches in vision-language models. Specifically, we compare two strategies in combination with sub-unit alignment. First, EEG and text representations may be projected using neural networks to a new, shared latent space prior to alignment (as in CLIP; [13, 12]), denoted henceforth as ELMe,l. Alternatively, the EEG embeddings may be projected into the output space of the language model (as in M-FLAG by [17]), denoted as ELMl.
@@
Dropping the recording subscript $i$ for brevity, each pair $\left(\mathbf{x}_{e},\mathbf{x}_{l}\right)$ is projected into the vectors $\mathbf{e}\in\mathbb{R}^{d}$ and $\mathbf{l}\in\mathbb{R}^{d}$ respectively. For every $\mathbf{x}_{e}$, text of the associated report is sampled according to $\tilde{\mathbf{x}}_{l}=z_{l}\left(\mathbf{x}_{l}\right)$, where $z_{l}$ represents the language sampling function detailed below. First, both the EEG crop $\mathbf{x}_{e}$ and text $\tilde{\mathbf{x}}_{l}$ are encoded into vectors $\mathbf{h}_{e}$ and $\mathbf{h}_{l}$. For ELMe,l, we use projectors $g_{e}$ and $g_{l}$ to yield vectors $\mathbf{e}$ and $\mathbf{l}$, whereas for ELMl the text embedding are not projected:
@@
As paired medical EEG data and clinical reports are scarce, training the text encoder function $f_{l}$ from scratch is unlikely to be successful. Furthermore, employing an existing language model and finetuning the model during multimodal pretraining can lead to training instability and collapse of the latent space [16, 17]. To prevent resulting information loss, we follow the recommendations by Liu et al. [17] to use a pretrained language model for $f_{l}$ and freeze its weights during training. We adopt their proposed composite loss for ELMl to learn $f_{e}$ and $g_{e}$:
@@
$\mathcal{L}_{orth}$ $=\sum_{j=1}\left(1-\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)_{jj}\right)^{2}+\sum_{j\neq k}\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)^{2}_{jk},$ (5)
@@
Meanwhile, ELMe,l relies on the cosine similarities between normalized EEG and text embeddings, $s^{e2l}_{j,j}=\mathbf{\hat{e}}_{j}^{\top}\mathbf{\hat{l}}_{j}$, and between text and EEG, $s^{l2e}_{j,j}=\mathbf{\hat{l}}_{j}^{\top}\mathbf{\hat{e}}_{j}$, with $j=1,2,3,...,N$ for batch size $N$ [12]. The multimodal contrastive loss uses a temperature hyperparameter $\tau$ (set to 0.3; Appendix A.4) and is formulated as:
$\mathcal{L}_{j,k}^{e2l}$ $=-\log\frac{\exp\left(s^{e2l}_{j,k}/\tau\right)}{\sum_{m=1}^{N}\exp\left(s^{e2l}_{j,m}/\tau\right)}$ (6)
$\mathcal{L}_{j,k}^{l2e}$ $=-\log\frac{\exp\left(s^{l2e}_{j,k}/\tau\right)}{\sum_{m=1}^{N}\exp\left(s^{l2e}_{j,m}/\tau\right)}$ (7)
$\mathcal{L}_{align}$ $=\frac{1}{2N}\sum_{j=1}^{N}\sum_{k=1}^{N}\left(\mathcal{L}_{j,k}^{e2l}+\mathcal{L}_{j,k}^{l2e}\right)$ (8)
Language encoder. For $f_{l}$ we use a transformer model which was pretrained in a contrastive manner on PubMed search logs (MedCPT [39]), which is based on PubMedBERT [40]. See Appendix A.3 for a comparison of language models. ELMl adopts the language model’s native hidden dimensionality (768), while for ELMe,l we project to a dimensionality of 256.
EEG encoder. For the EEG encoder $f_{e}$ we use a randomly initialized residual convolutional neural network, with an identical architecture across the compared training contexts: self-supervised learning with only EEG data and EEG-language pretraining. We use nonlinear MLPs with a single-hidden layer for $g_{e}$ and $g_{l}$, as well as for the projector head in EEG-only self-supervised learning. More details are provided in Appendix A.2.
@@ 3.1.2 EEG-only self-supervised learning
We compare the representations learned by EEG-language pretraining to those obtained via EEG-only pretraining. First, we employ multiple methods that train for invariance to data augmentations. This is achieved by sampling data augmentations for each EEG crop $\textbf{x}_{e}$, resulting in two differing data views $\{\textbf{x}_{e}^{\prime},\textbf{x}_{e}^{\prime\prime}\}$. It is important the data augmentations do not destroy the semantic information in $\textbf{x}_{e}$. Training to align the embeddings of these views while preventing collapse has been shown to yield data representations useful for downstream tasks [7, 8, 9]. We implement the following methods (Appendix A.5): Bootstrap-Your-Own-Latent (BYOL; [41]), Variance-Invariance-Covariance Regularization (VICReg; [42]), and Contrast with the World Representation (ContraWR; [9]). Additionally, we compare against methods using the temporal ordering of EEG crops: Relative Positioning (RP; [10]), Temporal Shuffling (TS; [10]), Contrastive Predictive Coding (CPC; [10]).
3.2 Datasets and preprocessing
• TUEG. The Temple University Hospital (TUH) EEG Corpus is the largest available corpus of hospital EEG data with varying montages, channel counts, and sampling frequencies (n=26846 [15]). For each patient, one or more EEG sessions are provided, each of which contains one or more recordings. For most of the dataset, no labels are available beyond patient age and sex. However, each EEG session is associated with a natural-language clinical report, containing information about the patient such as demographics and clinical history, as well as the physician’s impression of EEG recording and clinical status.
• TUAB. The TUH Abnormal EEG corpus is a subset of TUEG which was manually labeled by clinicians indicating whether the EEG displays pathological abnormalities [43]. This enables the binary classification task of predicting the status of {normal, abnormal}. As training (n=2717) and evaluation (n=276) sets are provided, we use the latter as a hold-out test set.

### v3

@@ Abstract
Multimodal language modeling has enabled breakthroughs for representation learning, yet remains unexplored in the realm of functional brain data for pathology detection. This paper pioneers EEG-language models (ELMs) trained on clinical reports and 15000 EEGs. We propose to combine multimodal alignment in this novel domain with timeseries cropping and text segmentation, enabling an extension based on multiple instance learning to alleviate misalignment between irrelevant EEG or text segments. Our multimodal models significantly improve pathology detection compared to EEG-only models across four evaluations and for the first time enable zero-shot classification as well as retrieval of both neural signals and reports. In sum, these results highlight the potential of ELMs, representing significant progress for clinical applications.
@@ 1 Introduction
Medical neuroimaging, including electroencephalography (EEG), has lagged behind other fields in leveraging the significant advancements of deep learning. While EEG sees widespread clinical use for pathology detection, in particular for epilepsy [1, 2] as well as sleep disorders [3], available annotated data is scarce. As the impressive scaling properties of deep learning are now well described [4, 5], self-supervised learning (SSL) is a promising direction by enabling pretraining with unlabeled data and thereby increasing available training sample sizes [6, 7]. Various such approaches have shown initial success when applied to EEG. These include methods relying on data-augmentations [8, 9], the temporal ordering of EEG data [10], as well as masking and reconstruction [11]. However, these are hindered by the difficulty of creating appropriate data augmentations and, especially reconstruction techniques, by low signal-to-noise. Thus, progress in the medical context has lagged, likely further exacerbated by the modality displaying high similarity between pathologies.
@@
Figure 1: Overview of the methodology. (Left) The ELM-MIL approach allows flexible multimodal alignment by cropping EEG and segmenting medical reports. We sample multiple positives in a cross-modal fashion, such that each EEG crop can be aligned with any number of segments from the paired report ($l|e$). Vice versa, text can be aligned selectively to crops across the EEG recording ($e|l$), illustrated by the differently shaded arrows. (Middle) An overview of investigated methods by visualizing the cross-modal similarity matrices. (Right) To evaluate models, we perform bidirectional retrieval analyses and perform multiple pathology detection tasks.
However, language-EEG pretraining also entails unique challenges. First, datasets are generally smaller than those used in radiology and especially computer vision. Second, the clinical reports tend to be highly heterogeneous. While previous applications have paired natural and medical images with short captions [12, 13], EEG reports tend to span multiple paragraphs and include information irrelevant to downstream clinical tasks, potentially hindering the pretraining process. Moreover, they do not contain temporal information about when events occurred during the recording.
The current work presents the application of aligning functional brain data with medical textual information for the first time by training EEG-language models (ELMs). To overcome the challenging formats of modalities, constituting long timeseries and multiparagraph reports, we propose sub-unit alignment. To address inconsistent relevance of EEG-text pairs, we additionally propose an extension drawing on insights from the field of multiple instance learning (MIL). Furthermore, we investigate how to best handle the heterogeneity of medical EEG reports. Specifically, we filter reports and perform content-based text segmentation, enabling inference on the relative importance of the different information sources. Our approach allows us to provide the first evidence of considerable retrieval capabilities for clinical reports and EEG. We furthermore test downstream performance of ELMs on classifying pathological EEG, which is a widespread clinical task, using four evaluations. These tests include zero-shot classification, leveraging language capabilities to show our approach’s flexibility. Our results constitute considerable increases in pathology detection performance, especially in scenarios with few labels. These are particularly relevant for clinical contexts, which tend to operate with smaller datasets compared to many common areas of deep learning applications.
@@ 2 Related work
  • Self-supervised learning with EEG data. SSL with EEG data has been predominantly applied to emotion recognition [16, 17], motor imagery [18, 19], sleep staging [9, 19], as well as pathology detection. For the latter application, the temporal order of EEG crops was used initially to demonstrate label-efficient representation learning [10]. Augmentation-based contrastive learning, combined with larger EEG encoders trained on multiple datasets, further improved pathology detection [8]. Recent studies have explored the use of transformers [20, 11], with a focus on scaling while adopting tokenization in an attempt to improve the challenge of effective cross-dataset EEG training.
  • Using EEG for pathology detection. While SSL shows good performance for pathology detection, it is particularly in contexts with little annotated data that it performs well. When more labeled data is available, expert-based feature extraction combined with traditional machine learning classifiers are competitive together with supervised deep learning [21, 22, 23, 24, 25]. This trend has also been observed in other EEG applications [26, 27]. This may indicate that inter-rater variability in EEG classification may create a performance ceiling [28, 22], highlighting the importance of improving classification with limited labels.
  • Medical multimodal language modeling. Medical vision-language modeling aims to guide self-supervised pretraining on medical images using textual information in reports, with performance on a variety of downstream tasks benefiting as a result [29, 30, 13]. Due to less available data in the medical domain, using a pretrained, frozen language encoder was found to boost downstream performance while considerably reducing computational cost [31]. Nevertheless, this line of work has focused mainly on the ECG, X-ray, CT images, and structural MRI images [32, 33, 34].
  • Multiple instance learning. MIL has seen only limited exploration for EEG. Initial studies have investigated the framework by casting crops of EEG as instances and training classifiers for emotion recognition [35], motor imagery [36], mental disorders [37], and sleep apnea [37]. Of these, only the latter has relied on deep learning.
@@ 3.1.1 EEG-language pretraining
Here we detail the setup for pretraining ELMs. Whereas vision-language models are typically trained by aligning a 2D image with a short caption [12, 13], EEG-language modeling is confronted with long EEG time series and multi-paragraph medical reports. To overcome this, we employ text segmentation and time series cropping to create multiple non-overlapping samples per modality and subject. Next, we propose sub-unit alignment by pretraining on these cropped samples. In addition to considerably increasing sample size, this enables the extension of successful approaches in vision-language models. We initially describe two strategies for sub-unit alignment. First, EEG and text representations may be projected using neural networks to a new, shared latent space prior to alignment (as in CLIP; [12, 13]), denoted henceforth as ELMe,l. Alternatively, the EEG embeddings may be projected into the output space of the language model (as in M-FLAG by [31]), denoted as ELMl and trained using a bespoke loss function. This approach was found to reduce latent collapse in smaller data settings [31]. Following a description of these models, we will introduce an extension based on MIL.
@@
Dropping the recording subscript $i$ for brevity, each pair $\left(\mathbf{x}_{e},\mathbf{x}_{l}\right)$ is projected into the vectors $\mathbf{e}\in\mathbb{R}^{d}$ and $\mathbf{l}\in\mathbb{R}^{d}$ respectively. For every $\mathbf{x}_{e}$, text of the associated report is sampled according to $\tilde{\mathbf{x}}_{l}=z_{l}\left(\mathbf{x}_{l}\right)$, where $z_{l}$ represents the language sampling function detailed below. First, both the EEG crop $\mathbf{x}_{e}$ and text $\tilde{\mathbf{x}}_{l}$ are encoded into vectors $\mathbf{h}_{e}$ and $\mathbf{h}_{l}$. For ELMe,l, we use projectors $g_{e}$ and $g_{l}$ to yield vectors $\mathbf{e}$ and $\mathbf{l}$, whereas for ELMl the text embeddings are not projected:
@@
As paired medical EEG data and clinical reports are scarce, training the text encoder function $f_{l}$ from scratch is unlikely to be successful. Furthermore, employing an existing language model and finetuning the model during multimodal pretraining can lead to training instability and collapse of the latent space [38, 31]. To prevent resulting information loss, we follow the recommendations by [31] to use a pretrained language model for $f_{l}$ and freeze its weights during training. For ELMl, we adopt their proposed composite loss to learn $f_{e}$ and $g_{e}$:
@@
$\mathcal{L}_{orth}$ $=\sum_{j=1}\left(1-\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)_{jj}\right)^{2}$ (5)
$+\sum_{j\neq k}\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)^{2}_{jk}$ (6)
@@
Meanwhile, ELMe,l relies on the cosine similarities between normalized EEG and text embeddings, $s^{e2l}_{j,j}=\mathbf{\hat{e}}_{j}^{\top}\mathbf{\hat{l}}_{j}$, and between text and EEG, $s^{l2e}_{j,j}=\mathbf{\hat{l}}_{j}^{\top}\mathbf{\hat{e}}_{j}$, with $j=1,2,3,...,B$ for batch size $B$ [12]. The multimodal contrastive InfoNCE loss uses a temperature hyperparameter $\tau$ (set to 0.3 using a holdout set; Appendix B.4) and is formulated as:
$\mathcal{L}_{j,k}^{e2l}$ $=-\log\frac{\exp\left(s^{e2l}_{j,k}/\tau\right)}{\sum_{m=1}^{B}\exp\left(s^{e2l}_{j,m}/\tau\right)}$ (7)
$\mathcal{L}_{j,k}^{l2e}$ $=-\log\frac{\exp\left(s^{l2e}_{j,k}/\tau\right)}{\sum_{m=1}^{B}\exp\left(s^{l2e}_{j,m}/\tau\right)}$ (8)
$\mathcal{L}_{align}$ $=\frac{1}{2B}\sum_{j=1}^{B}\sum_{k=1}^{B}\left(\mathcal{L}_{j,k}^{e2l}+\mathcal{L}_{j,k}^{l2e}\right)$ (9)
Multiple instance learning. While previous approaches aim to align text and EEG crops uniformly, certain text segments likely describe specific EEG sections more accurately than others. Therefore, we introduce a MIL alignment strategy that builds on ELMe,l and accommodates multiple positive samples, allowing for more nuanced multimodal relationships. Whereas MIL approaches often rely on operations such as max-pooling to focus on single positive samples, we rely on insights from the video-text alignment approach (MIL-NCE) by [39]. For a given text sample $\mathbf{x}_{l}$, we sample multiple positive EEG crops $\mathbf{x}_{e}$ from the paired recording to approximate the $P(e|l)$ distribution, while for an EEG crop, multiple text segments are sampled to model the $P(l|e)$ distribution. We combine these and sample positives for each EEG and text segment respectively to approximate $P(e,l)$ via bidirectional alignment. This approach effectively relaxes the assumption of strong alignment for each individual $\left(\mathbf{x}_{e},\mathbf{x}_{l}\right)$ pair, instead assuming that, on average, positive samples should have higher similarity scores than negative samples. To this end, we extend the InfoNCE loss to multiple instances:
$\mathcal{L}^{e|l}$ $=-\frac{1}{B_{l}}\sum_{k=1}^{B_{l}}\log\frac{\frac{1}{|P_{k}|}\sum_{j\in P_{k}}\exp(s^{e2l}_{j,k}/\tau)}{\sum_{j=1}^{B_{e}}\exp(s^{e2l}_{j,k}/\tau)}$ (10)
$\mathcal{L}^{l|e}$ $=-\frac{1}{B_{e}}\sum_{k=1}^{B_{e}}\log\frac{\frac{1}{|Q_{k}|}\sum_{j\in Q_{k}}\exp(s^{l2e}_{j,k}/\tau)}{\sum_{j=1}^{B_{l}}\exp(s^{l2e}_{j,k}/\tau)}$ (11)
with |$\text{with }|P_{k}|\leq N\ \text{ and }|Q_{k}|\leq M,$ (12)
$\mathcal{L}^{e,l}$ $=\frac{1}{2}\left(\mathcal{L}^{e|l}+\mathcal{L}^{l|e}\right)$ (13)
and where $P_{k}$ and $Q_{k}$ are sets of positive EEG and text segments respectively. $B_{e}$ and $B_{l}$ are the batch sizes for EEG and text respectively, for which we sample up to $N$ EEG and $M$ text segments for $\frac{B}{N}$ subjects. We set $N=32$ and $M=8$ as this covers all samples for a majority of subjects. We normalize using $|Q_{k}|$ or $|P_{k}|$ to account for the varying number of crops across subjects.
Language encoder. For $f_{l}$ we use a transformer model which was pretrained with medical data in a contrastive manner on PubMed search logs (MedCPT; [40]). See Appendix B.3 for a comparison of language models. ELMl adopts the language model’s native hidden dimensionality (768), while for ELMe,l and ELM-MIL we project to a dimensionality of 256.
EEG encoder. For the EEG encoder $f_{e}$ we use a randomly initialized residual convolutional neural network, with an identical backbone architecture across all comparisons. We use nonlinear MLPs with a single-hidden layer for $g_{e}$ and $g_{l}$, as well as for the projector head in EEG-only self-supervised learning. More details are provided in Appendix B.2.
@@ 3.1.2 EEG-only self-supervised learning
We compare the representations learned by EEG-language pretraining to those obtained via EEG-only pretraining using the same pretraining dataset and EEG encoder. We provide further information on the following methods in Appendix B.5: Bootstrap-Your-Own-Latent (BYOL; [41]), Variance-Invariance-Covariance Regularization (VICReg; [42]), Contrast with the World Representation (ContraWR; [9]), Relative Positioning (RP; [10]), Temporal Shuffling (TS; [10]), Contrastive Predictive Coding (CPC; [10]).
3.2 Datasets and evaluation tasks
• TUEG. The Temple University Hospital (TUH) EEG Corpus is the largest available corpus of hospital EEG data with varying montages, channel counts, and sampling frequencies (n=26846 [15]). For most of the dataset, no labels are available beyond patient age and sex. However, many EEG sessions are associated with a natural-language clinical report.
• TUAB. The TUH Abnormal EEG corpus is a subset of TUEG which was manually labeled by clinicians indicating whether the EEG displays pathological abnormalities [43]. This enables the binary classification task of predicting the status of {normal, abnormal} on a recording-level. Following the literature, we use the provided evaluation set as the hold-out test set.
• NMT. We leverage the NMT Scalp EEG Dataset [44] in order to validate our results for recording-level abnormality classification out-of-distribution. The NMT dataset deviates considerably from TUEG. Data was recorded from a South Asian population in Pakistan, using a different EEG recording setup. Furthermore, the NMT participants are considerably younger, feature more males (66.6%), and their EEG recordings are labeled predominantly normal (83.8% in the training set, while the test set is balanced). We use the provided train/test split.
To further evaluate learned representation, we use tasks requiring classification of single, short 5-second EEG crops.
• TUSZ. The TUH Seizure Detection Corpus [45] is a subset of TUEG which has sections labeled to contain either seizure or background activity. We perform binary classification using 5-fold cross validation on the provided train and dev sets (n=6491), while testing on the eval set.
• TUEV. An Events Corpus [15] containing annotated EEG with six classes, of which three are clinical (spike and slow wave, generalized periodic epileptiform discharge, periodic lateralized epileptiform discharge) as well as eye movements, artifacts, and background activity. We only use the provided train set (5-fold CV) due to the test set not including the TUEG subject identifiers, which would have prevented the exclusion of these subjects from the pretraining data. For each 1-second event, we include two seconds of context before and after [11].
Table 1: Dataset sample sizes. Crop lengths are indicated for ELM/EEG-only models.
Data subset EEG files Reports Crop Length
TUEG Pretrain 15144 11785 Variable
TUAB train 2712 Not used 60/20s
TUAB test 276 Not used 60/20s
Retrieval test 437 437 60/-
NMT Train 2216 - 60/20s
NMT Test 183 - 60/20s
TUSZ Train+Dev 6491 - 5/5s
TUSZ Test 865 - 5/5s
TUEV Train 359 - 5/5s
Figure 2: A,B) A set of EEG-Language models are evaluated on their retrieval ability using top-k accuracy out of 437 patients. C) Linear probing performance on TUAB with 1% of labels and D) zero-shot classification. Error bars indicate standard deviations over five model training runs.
3.3 Preprocessing